OpenStudy (anonymous):

Ashley is comparing three investment accounts offering different rates. Option A: APR of 4.95% compounding monthly Option B: APR of 4.85% compounding quarterly Option C: APR of 4.895% compounding daily She would like to earn at least a 5% annual yield. Which account(s) will give Alexandra the yield she wants? Options A and B only Option B only Options A and C only Option C only

OpenStudy (tkhunny):

\(\left(1 + \dfrac{0.0495}{12}\right)^{12}\) \(\left(1 + \dfrac{0.0485}{4}\right)^{4}\) \(\left(1 + \dfrac{0.04895}{365}\right)^{365}\) Exactly the same sort of construction. Evaluate!